### v1

# A randomized small-block Lanczos method for large-scale null space computations

Abstract
Computing the null space of a large sparse matrix $A$ is a challenging computational problem, especially if the nullity—the dimension of the null space—is large. When using a block Lanczos method for this purpose, conventional wisdom suggests to use a block size $d$ that is not smaller than the nullity. In this work, we show how randomness can be utilized to allow for smaller $d$ without sacrificing convergence or reliability. Even $d=1$, corresponding to the standard single-vector Lanczos method, becomes a safe choice. This is achieved by using a small random diagonal perturbation, which moves the zero eigenvalues of $A^{\mathsf{T}}A$ away from each other, and a random initial guess. We analyze the effect of the perturbation on the attainable quality of the null space and derive convergence results that establish robust convergence for $d=1$. As demonstrated by our numerical experiments, a smaller block size combined with restarting and partial reorthogonalization results in reduced memory requirements and computational effort. It also allows for the incremental computation of the null space, without requiring a priori knowledge of the nullity.
keywords:
Sparse matrix, null space, Krylov subspace method, randomized algorithm
{AMS}
65F15, 65F50, 68W20
1 Introduction
This work is concerned with computing the null space of a large sparse matrix $A\in\mathbb{R}^{m\times n}$, a problem that finds numerous applications. For example, the connected components of an undirected graph can be identified by the null space of the graph Laplacian [18]. In computational topology, the null space of cohomology matrices reflects the generators of the cohomology groups. The nullity helps determine the Betti numbers of a topological space [7]. Furthermore, null space calculations play a crucial role in various engineering problems, such as system dimensionality reduction [34], finite element models [29], linear equality constrained state estimation [13], grasp control [25], computer vision [36], and multivariate polynomial root finding [2].
In several of these applications, the dimension of the null space is not necessarily small and can easily reach a few hundred or even thousands. For small- to medium-sized matrices, the singular value decomposition (SVD) [10] or a suitably modified QR decomposition [8] are the methods of choice for computing null spaces. For large-sized matrices, these dense methods become too expensive. There are several structured null space solvers to deal with specific large-scale situations, such as [3] for banded matrices, [11] for matrices with small nullity and amenable to sparse LU factorization with partial pivoting, [9] for sparse matrices amenable to a rank-revealing sparse QR factorization [4], and [22] for a very tall and skinny matrix.
When $A$ is large and has no particular structure beyond being sparse, one needs to resort to Krylov subspace methods, which utilize matrix-vector products with $A$ and its transpose $A^{\mathsf{T}}$. Let $N$ denote the null space dimension of $A$. When $N=1$, standard choices include the Lanczos method for computing the smallest eigenvalue and eigenvector of $A^{\mathsf{T}}A$ or the mathematically equivalent, but numerically more reliable Golub–Kahan bidiagonalization directly applied to $A$ [10, Chapter 10]. When $N>1$, the Lanczos method (with a single starting vector) becomes unreliable and prone to miss components of the null space. In fact, in exact arithmetic only a one-dimensional subspace of the null space can be extracted via Lanczos [23, Section 12.2]. Block Lanczos methods address this issue by utilizing a block of $d>1$ starting vectors instead of a single starting vector [23, Section 13.10]. Common wisdom (see [23, Section 13.12] for examples) and existing convergence analyses all suggest to choose $d$ not smaller than the size of the eigenvalue cluster of interest. Applied to null space computation this would require $d\geq N$ and in exact arithmetic this condition is indeed necessary. The presence of round-off error blurs the situation. As an example, consider the $420\times 420$ diagonal matrix
(1) $A=\mathrm{diag}(0,\dotsc,0,1,2,\dotsc,399),$
that is, the first $21$ diagonal entries are zero. When applying block Lanczos to $A$ in order to compute its null space and choosing a block of $d$ (Gaussian) random starting vectors, one would expect to obtain no more than $d$ (approximate) zero eigenvalues and corresponding eigenvectors. However, when executing block Lanczos in double precision arithmetic without breakdown and declaring eigenvalues less than $10^{-4}$ as zero, we obtained the results reported in Table 1.
Table 1: Null space dimension obtained when applying block Lanczos with $d$ random starting vectors to the matrix from Eq. 1.
Block size $d$ 1 2 3 4 5 6 7 10 12 14 15 20 21
Dimension 11 11 16 12 15 15 21 20 21 21 21 21 21
While the null space dimension estimated by block Lanczos for small values of $d$ is erratic and smaller than $21$, it is always larger than $d$. In fact, already for $d=12$, the correct null space dimension is detected. Experiments with other matrices lead to similar findings, suggesting that round-off error breaks some of the curse incurred by eigenvalue clusters, but it does not fully address the convergence issues of single-vector or small-block Lanczos methods either.
The situation described above is reminiscent of recent work by Meyer, Musco and Musco [19] on the single-vector Lanczos method for low-rank approximation. Taking the randomness of the starting vector into account they establish a convergence result that still requires the (larger) singular values to be distinct but the dependence of the complexity on the gaps between eigenvalues is very mild, in fact logarithmic. Although single-vector or small-block Krylov subspace methods do not satisfy gap-independent convergence bounds [16, 27, 21, 35], the theory in [19] suggests that small random perturbations of $A$ can easily break the adverse role of repeated singular values. As we have seen, perturbations due to round-off are not sufficient to achieve this effect, but (slightly larger) random perturbations will do, thanks to eigenvalue repulsion results.
Inspired by [19], we propose a randomized small-block Lanczos method for null space calculation, outlined in Algorithm 1. Compared to the task of low-rank approximation, there are two major differences. First, in low-rank approximation, the largest few singular values are usually unknown and, in most applications, they do not form tight large clusters. In the context of null spaces, the desired singular values form one large cluster of zeros. Second, while convergence to the relevant invariant subspace is not necessary for low-rank approximation [6], such convergence is imperative to obtain a good null space approximation.
Input: Matrix $A\in\mathbb{R}^{m\times n}$ with $m\geq n$ and $N$ zero singular values. Perturbation parameter $\epsilon>0$. Block size $d$. Number of iterations $\ell$.
Output: Orthonormal basis $V$ approximating null space of $A$.
1 Set $B=A^{\mathsf{T}}A+\epsilon D$, with a diagonal matrix $D$ having diagonal entries uniformly i.i.d. in $[0,1]$;
2 Draw Gaussian random matrix $\Omega\in\mathbb{R}^{n\times d}$;
3 Perform block Lanczos to compute Lanczos decomposition $BZ_{\ell}=Z_{\ell}T_{\ell}+Q_{\ell+1}E_{\ell+1}$, where $Z_{\ell}$ is an orthonormal basis of Krylov subspace $\mathrm{span}\{\Omega,B\Omega,\dotsc,B^{\ell-1}\Omega\}$ and $T_{\ell}$ is block tridiagonal;
4 Compute an orthonormal basis $V_{Z}$ for the invariant space associated with $N$ smallest eigenvalues of $T_{\ell}$;
5 return $V=Z_{\ell}V_{Z}$;
Algorithm 1 Randomized small-block Lanczos for null space computation
Remark 1.1.
Line 1 of Algorithm 1 extracts the null space approximation from Ritz vectors by computing the spectral decomposition of $T_{\ell}=Z_{\ell}^{\mathsf{T}}BZ_{\ell}$. Alternatively, one could remove the perturbation and compute the spectral decomposition of $Z_{\ell}^{\mathsf{T}}A^{\mathsf{T}}AZ_{\ell}$ or, equivalently, the SVD of $AZ_{\ell}$. Preliminary numerical experiments indicate that this only yields minor accuracy improvements, at best, not justifying the extra computational effort needed to perform the alternative procedure.
The rest of this paper is organized as follows.
In Section 2, we analyze the impact of the random perturbation $\epsilon D$ introduced by Algorithm 1. In particular, Theorem 2.2 provides a new eigenvalue repulsion result for the perturbed zero eigenvalues of $A^{\mathsf{T}}A$. Our result differs from an existing result [19, Lemma 12] in two aspects: (1) By focusing on the repulsion of zero eigenvalues rather than the entire spectrum, we establish $\mathcal{O}(\epsilon)$ gaps instead of $\mathcal{O}(\epsilon^{2})$ gaps. (2) By considering relative eigenvalue gaps, we establish lower bound independent of the norm of $A$. At the same time, the perturbation incurs a limit on the attainable accuracy of the null space approximation, and we quantify this effect by a perturbation analysis.
Section 3 performs a self-contained convergence analysis of the randomized single-vector Lanczos method (corresponding to $d=1$ in Algorithm 1). While our analysis shares similarities with [19], we achieve more refined results by focusing on $d=1$ and considering the convergence of individual Ritz vectors. This analysis cumulates in Theorems 3.1 and 3.10, establishing bounds on the null space approximation, whose sharpness is verified by the numerical results.
In Section 4, we consider several practical improvements of Algorithm 1, including preconditioning, restarting, and partial reorthogonalization.
Finally, Section 5 contains several numerical results for applications, such as the computation of connected graph components and cohomology computation. These numerical experiments not only validate our theoretical findings regarding the randomized single-vector Lanczos method but also highlight the efficiency of the small-block Lanczos method. Notably, it highlights when our method is preferable compared to other null space solvers, particularly when the nullity is substantial and memory constraints are a limiting factor.
Assumptions and notation
Throughout this paper we assume that $A\in\mathbb{R}^{m\times n}$ satisfies $m\geq n$ and has $N\geq 1$ zero singular values. In accordance with Step 1 of Algorithm 1, we define
(2) $B:=A^{\mathsf{T}}A+\epsilon D,$
where $D$ is diagonal with nonnegative diagonal entries and $\epsilon>0$ controls the strength of the perturbation. We let $\lambda_{1},\ldots,\lambda_{n}$ denote the (nonnegative) eigenvalues of $B$, sorted such that
$0\leq\lambda_{1}\leq\dotsb\leq\lambda_{n}.$
[…]
Remark 1.2.
For null space computation, the cases $m\geq n$ and $m<n$ are fundamentally different because $m<n$ implies that $A$ has a null space of dimension at least $n-m$. When $n$ is much larger than $m$, say $n=2m$, one needs to resort to dense methods like the SVD, in the absence of additional structure.
2 Perturbation analysis
The purpose of this section is to quantify the impact of the (random) perturbation $\epsilon D$ introduced by Algorithm 1.
2.1 Perturbation and repulsion of zero eigenvalues
We start by considering the impact of the perturbation on the $N$ zero eigenvalues of $A^{\mathsf{T}}A$. Suppose that $0<\epsilon\|D\|<\underline{\sigma}^{2}$, where $\underline{\sigma}$ denotes the smallest nonzero singular value of $A$. Then the zero and nonzero eigenvalues of $A^{\mathsf{T}}A$ will not be mixed by the perturbation $\epsilon D$. Indeed, using that $D$ is positive semi-definite, it follows from Weyl’s inequality [14, Theorem 4.3.1] that the eigenvalues $\lambda_{1},\dotsc,\lambda_{n}$ of $B$ satisfy
(3) $0\leq\lambda_{1}\leq\dotsb\leq\lambda_{N}\leq\epsilon\|D\|<\underline{\sigma}^{2}\leq\lambda_{N+1}\leq\dotsb\leq\lambda_{n}.$
Theorem 2.2 below shows that it is unlikely that the perturbed zero eigenvalues $\lambda_{1},\dotsc,\lambda_{N}$ remain very close to each other when $D$ is randomly chosen. To establish this eigenvalue repulsion result, we follow the proof of [19, Lemma 12] and make use of the following proposition.
Proposition 2.1 (Equation (1.11) in [1] attributed to [20]).
Let $B_{0}\in\mathbb{R}^{n\times n}$ be symmetric, and let $D\in\mathbb{R}^{n\times n}$ be a random diagonal matrix with the diagonal entries drawn i.i.d. from a distribution with bounded probability density function $\rho(\cdot)$. Then for any bounded interval $\mathcal{I}\subset\mathbb{R}$ and $\epsilon>0$, the inequality
$\Pr(B_{0}+\epsilon D\text{ has at least $2$ eigenvalues in }\mathcal{I})\leq C_{0}(\epsilon^{-1}\lVert\rho\rVert_{\infty}\lvert\mathcal{I}\rvert n)^{2},$
holds for some universal constant $C_{0}$, where $\lvert\mathcal{I}\rvert$ is the length of $\mathcal{I}$.
Theorem 2.2.
Consider the matrix $B$ defined in (2) and suppose that the diagonal entries of $D$ are drawn i.i.d. from the uniform distribution on $[0,1]$. Let $\underline{\sigma}$ denote the smallest nonzero singular value of $A$. If $0<\epsilon<\underline{\sigma}^{2}$, the inequality
$\min_{1\leq i\leq N-1}\lvert\lambda_{i+1}-\lambda_{i}\rvert\geq\frac{\delta\epsilon}{Cn^{2}},$
holds with probability at least $1-\delta$ for any $10<\delta<$ 1, where $C$ is a universal constant.
Proof 2.3.
Because of $\|D\|\leq 1$, it follows from (3) that $\lambda_{1},\dotsc,\lambda_{N}$ are contained in the interval $[0,\epsilon]$. We cover this interval with overlapping (smaller) intervals $\mathcal{I}_{j}=[j\gamma,(j+2)\gamma]$ for $j=0,\ldots,\epsilon/\gamma-2$, where—for the moment—$\gamma$ is an arbitrary number such that $20<\gamma\leq\epsilon/$ 2 and $\epsilon/\gamma$ is an integer. If two neighboring eigenvalues satisfy $\lambda_{i+1}-\lambda_{i}\leq\gamma$, then there clearly exists an interval $\mathcal{I}_{j}$ such that $\lambda_{i},\lambda_{i+1}\in\mathcal{I}_{j}$. Thus,
$\Pr\{\exists i\colon\lambda_{i+1}-\lambda_{i}\leq\gamma\}$$\leq$ $\Pr\{\exists i,\,j\colon\lambda_{i},\lambda_{i+1}\in\mathcal{I}_{j}\}\leq\sum_{j=0}^{\epsilon/\gamma-2}\Pr\{\exists i\colon\lambda_{i},\lambda_{i+1}\in\mathcal{I}_{j}\}$
$\leq$ $\sum_{j=0}^{\epsilon/\gamma-2}4C_{0}n^{2}\epsilon^{-2}\gamma^{2}=4C_{0}n^{2}\gamma\epsilon^{-1}\big{(}1-\gamma\epsilon^{-1}\big{)}\leq\frac{4C_{0}n^{2}}{\epsilon/\gamma+1},$
where we used the union bound and Proposition 2.1 in the second and third inequalities, respectively. The last inequality uses $x(1-x)=x(1-x^{2})/(1+x)\leq x/(1+x)$ with $x=\gamma\epsilon^{-1}$. The proof is concluded by setting $C=4C_{0}$, $\gamma=\epsilon/\lfloor\frac{Cn^{2}}{\delta}\rfloor\geq\frac{\delta\epsilon}{Cn^{2}}$ and using ${\epsilon}/{\gamma}=\lfloor{Cn^{2}}/{\delta}\rfloor>{Cn^{2}}/{\delta}-1$.
2.2 Perturbation of null space residual
We will bound the impact of perturbations on the null space residual of $A$ using the following eigenvalue perturbation result.
Proposition 2.4.
Consider symmetric matrices $B_{0}$ and $B\in\mathbb{R}^{n\times n}$ with eigenvalues $0\leq\lambda_{1}^{(0)}\leq\dotsb\leq\lambda^{(0)}_{n}$ and $0\leq\lambda_{1}\leq\dotsb\leq\lambda_{n}$, respectively. Let $V_{N}$ be an orthonormal basis for the invariant subspace of $B$ associated with $\lambda_{1},\dotsc,\lambda_{N}$. Suppose that $B-B_{0}$ is positive semi-definite, then
$\lVert V_{N}^{\mathsf{T}}B_{0}V_{N}\rVert\leq\lambda_{N}^{(0)}+\lVert B-B_{0}\rVert.$
If, additionally, $\lVert B-B_{0}\rVert<\lambda_{N+1}^{(0)}-\lambda_{N}^{(0)}$ then
$\lVert V_{N}^{\mathsf{T}}B_{0}V_{N}\rVert\leq\lambda_{N}^{(0)}+\frac{\lVert B-B_{0}\rVert^{2}}{\lambda_{N+1}^{(0)}-\lambda_{N}^{(0)}-\lVert B-B_{0}\rVert}.$
Proof 2.5.
The first result is a direct consequence of Weyl’s inequality:
$\lVert V_{N}^{\mathsf{T}}B_{0}V_{N}\rVert\leq\lVert V_{N}^{\mathsf{T}}BV_{N}\rVert=\lambda_{N}\leq\lambda_{N}^{(0)}+\lVert B-B_{0}\rVert,$
The second result follows from a quadratic residual bound for eigenvalues in [17, Theorem 1]. To apply this bound, we let $V_{\perp}$ be such that $[V_{N},V_{\perp}]$ is an $n\times n$ orthogonal matrix and consider the gap
(4) $g$ $=\lambda_{\min}\big{(}V_{\perp}^{\mathsf{T}}B_{0}V_{\perp}\big{)}-\lambda_{N}^{(0)}\geq\lambda_{\min}\big{(}V_{\perp}^{\mathsf{T}}BV_{\perp}\big{)}-\lambda_{N}^{(0)}-\|B-B_{0}\|$
$=\lambda_{N+1}-\lambda_{N}^{(0)}-\|B-B_{0}\|\geq\lambda_{N+1}^{(0)}-\lambda_{N}^{(0)}-\|B-B_{0}\|>0.$
Now, [17, Theorem 1] yields
$\lVert V_{N}^{\mathsf{T}}B_{0}V_{N}\rVert-\lambda_{N}^{(0)}\leq g^{-1}\lVert(I-V_{N}V_{N}^{\mathsf{T}})B_{0}V_{N}\rVert^{2}\leq g^{-1}\lVert B-B_{0}\rVert^{2},$
where the second inequality uses that $V_{N}$ spans an invariant subspace of $B$. Combined with Eq. 4, this proves the desired result.
Proposition 2.4 provides two kinds of perturbation bounds. The first bound does not require a gap but depends linearly on $\lVert B-B_{0}\rVert$, while the second one improves this to quadratic dependence at the expense of requiring a gap between $\lambda_{N}^{(0)}$ and $\lambda_{N+1}^{(0)}$.
Theorem 2.6.
Let $V_{N}\in\mathbb{R}^{n\times N}$ be an orthonormal basis of the invariant subspace associated with the $N$ smallest eigenvalues of $B=A^{\mathsf{T}}A+\epsilon D$ for a symmetric positive semi-definite matrix $D$ with $\lVert D\rVert\leq 1$. If $20<\epsilon<\underline{\sigma}^{2}/$ 2, where $\underline{\sigma}$ denotes the smallest positive singular value of $A$, then
$\lVert AV_{N}\rVert\leq\epsilon(\underline{\sigma}^{2}-\epsilon)^{-1/2}\leq\sqrt{\epsilon}.$
Proof 2.7.
The result essentially follows from inserting $B_{0}=A^{\mathsf{T}}A$ into the second bound of Proposition 2.4.
2.3 Principal angles and null space residual
To study the impact of perturbations on subspaces, we recall the definition and basic properties of principal angles. For later purposes, it will be helpful to admit subspaces of unequal dimension, as in [10, Section 6.4.3] and [38].
Definition 2.8.
Let the columns of $V_{0}\in\mathbb{R}^{n\times k_{0}}$ and $V_{1}\in\mathbb{R}^{n\times k_{1}}$ form orthonormal bases of subspaces $\mathcal{V}_{0}$ and $\mathcal{V}_{1}$, respectively. Let $k=\min\{k_{0},k_{1}\}$ and let $1\geq\sigma_{1}\geq\sigma_{2}\geq\dotsb\geq\sigma_{k}\geq 0$ denote the singular values of $V_{0}^{\mathsf{T}}V_{1}$. Then the principal angles $\angle_{t}(V_{0},V_{1})\in[0,\pi/2]$, $t=1,\dotsc,k$, between $\mathcal{V}_{0}$ and $\mathcal{V}_{1}$ are defined by
$\cos\angle_{t}(\mathcal{V}_{0},\mathcal{V}_{1})=\sigma_{t}.$
For brevity, we will write $\angle_{t}(\Omega_{0},\Omega_{1})$ instead of $\angle_{t}(\mathcal{V}_{0},\mathcal{V}_{1})$ for arbitrary, not necessarily orthonormal, basis for $\mathcal{V}_{0}$ and $\mathcal{V}_{1}$.
Note that the principal angles in Definition 2.8 are sorted in increasing order, and it is often the largest principal angle $\angle_{k}(V_{0},V_{1})$ that is of interest. In particular, it is well known that $\sin\angle_{k}(V_{0},V_{1})=\lVert V_{0}V_{0}^{\mathsf{T}}-V_{1}V_{1}^{\mathsf{T}}\rVert$ holds when $k_{0}=k_{1}$, implying that the orthogonal projectors are close for small $\angle_{k}(V_{0},V_{1})$. The other principal angles give partial information; knowing that some of them are small allows us to conclude that $\mathcal{V}_{0}$ and $\mathcal{V}_{1}$ nearly share a common subspace. In the context of null spaces, the following result establishes a residual bound in terms of principal angles.
Proposition 2.9.
Consider the setting of Definition 2.8 and a matrix $A\in\mathbb{R}^{m\times n}$. Then for every $t=1,\dotsc,k$ there exists a matrix $W_{t}\in\mathbb{R}^{k_{1}\times t}$ with orthonormal columns such that
$\lVert AV_{1}W_{t}\rVert\leq\lVert AV_{0}\rVert+\lVert A\rVert\sin\angle_{t}(V_{0},V_{1}).$
Proof 2.10.
Let $W_{t}$ be the right singular vectors corresponding to the largest $t$ singular values of $V_{0}^{\mathsf{T}}V_{1}$. Then
$\|(I-V_{0}V_{0}^{\mathsf{T}})V_{1}W_{t}\|^{2}$ $=$ $\max_{\|x\|=1}\|(I-V_{0}V_{0}^{\mathsf{T}})V_{1}W_{t}x\|^{2}$
$=$ $\max_{\|x\|=1}\big{(}\|V_{1}W_{t}x\|^{2}-\|V_{0}V_{0}^{\mathsf{T}}V_{1}W_{t}x\|^{2}\big{)}$
$=$ $1-\min_{\|x\|=1}\|V_{0}^{\mathsf{T}}V_{1}W_{t}x\|^{2}=1-\sigma_{t}^{2}=\sin^{2}\angle_{t}(V_{0},V_{1}).$
The proof is concluded by the triangular inequality:
$\lVert AV_{1}W_{t}\rVert$ $\leq\lVert AV_{0}V_{0}^{\mathsf{T}}V_{1}W_{t}\rVert+\lVert A(I-V_{0}V_{0}^{\mathsf{T}})V_{1}W_{t}\rVert$
$\leq\lVert AV_{0}\rVert+\lVert A\rVert\sin\angle_{t}(V_{0},V_{1}).$
We conclude this section with a useful characterization of the tangents of principal angles from [38, Theorem 3.1].
Proposition 2.11.
Let $[V_{0},V_{\perp}]$ be an orthonormal matrix, where $V_{0}\in\mathbb{R}^{n\times k_{0}}$. Suppose $\Omega_{1}\in\mathbb{R}^{n\times k_{1}}$ has column full rank. If $\mathrm{rank}(V_{0}^{\mathsf{T}}\Omega_{1})=\mathrm{rank}(\Omega_{1})\leq k_{0}$, then
$\lVert V_{\perp}^{\mathsf{T}}\Omega_{1}(V_{0}^{\mathsf{T}}\Omega_{1})^{\dagger}\rVert=\tan\angle_{k_{1}}(V_{0},\Omega_{1}).$
3 Convergence analysis
In this section, we study the convergence of Algorithm 1, focusing on the case of a single random starting vector $\omega$, that is, $d=1$.
3.1 Convergence of single-vector Krylov subspaces for small eigenvalues
In the following, we analyze how well the Krylov subspace generated in line 3 of Algorithm 1 contains approximations to (portions of) the invariant subspace associated with the $N$ smallest eigenvalues of $B$. The analysis mimics analogous results obtained in [19] for low-rank approximation. Compared to classical results [28, Chapter VI], our analysis differs by considering the convergence to a whole subspace and leveraging the randomness of the initial vector. In the following theorem, the main result of this section; $\mathsf{Cheb}_{k}(\lambda)$ denotes the Chebyshev polynomial of degree $k$ defined as
(5) $\mathsf{Cheb}_{k}(\lambda)=\frac{1}{2}\big{(}\big{(}\lambda+\sqrt{\lambda^{2}-1}\big{)}^{k}+\big{(}\lambda+\sqrt{\lambda^{2}-1}\big{)}^{-k}\big{)}.$
Theorem 3.1.
For a symmetric matrix $B\in\mathbb{R}^{n\times n}$ with eigenvalues $\lambda_{1}<\lambda_{2}<\dotsb<\lambda_{N}<\lambda_{N+1}\leq\dotsb\leq\lambda_{n}$, let $V_{N}\in\mathbb{R}^{n\times N}$ be an orthonormal basis of the invariant subspace associated with $\lambda_{1},\dotsc,\lambda_{N}$. Choose integers $t,\ell$ such that $1\leq t\leq N$ and $t\leq\ell\leq n$. Let $K_{\ell}=[\omega,B\omega,\dotsc,B^{\ell-1}\omega]$, where $\omega\in\mathbb{R}^{n}$ is a Gaussian random vector. Then, for any $10<\delta<$ 1, the inequality
(6) $\tan\angle_{t}(V_{N},K_{\ell})\leq\frac{N\sqrt{5\pi(n-N)\log(n/\delta)}}{\sqrt{2}\delta(t-1)!\cdot\mathsf{Cheb}_{\ell-t}\bigl{(}1+2\frac{\lambda_{N+1}-\lambda_{N}}{\lambda_{n}-\lambda_{N+1}}\bigr{)}}\Bigl{(}\frac{2}{\mathsf{relgap}}\Bigr{)}^{t-1}$
holds with probability at least $1-2\delta$, where
(7) $\mathsf{relgap}:=\min_{1\leq i\leq N-1}(\lambda_{i+1}-\lambda_{i})/(\lambda_{n}-\lambda_{1}).$
Combining Theorem 3.1 with Theorem 2.2, we obtain the following result.
Corollary 3.2.
Suppose that the matrix $B$ in Theorem 3.1 takes the form $B=A^{\mathsf{T}}A+\epsilon D$, where $A$ is an $m\times n$ matrix with null space dimension $N$ and $D$ is a random diagonal matrix with the diagonal entries drawn i.i.d. from the uniform distribution on $[0,1]$. Let $\underline{\sigma}$ and $\overline{\sigma}$ be the smallest and largest nonzero singular values of $A$, respectively. If $0<\epsilon<\underline{\sigma}^{2}$, then
(8) $\tan\angle_{t}(V_{N},K_{\ell})\leq\frac{N\sqrt{(n-N)\log(n/\delta)}\bigl{(}C(n^{2}\overline{\sigma}^{2})/(\delta\epsilon)\bigr{)}^{t-1}}{\delta\cdot(t-1)!\cdot\mathsf{Cheb}_{\ell-t}\bigl{(}1+\frac{2(\underline{\sigma}^{2}-\epsilon)}{\overline{\sigma}^{2}-\underline{\sigma}^{2}+\epsilon}\bigr{)}}$
holds with probability at least $1-3\delta$, where $C$ is a universal constant.
The rest of this section is concerned with the proof of Theorem 3.1, which follows the ideas from [19] and is divided into two parts. First, in Section 3.1.1, we will establish Eq. 6 for $\ell=t$, that is,
(9) $\tan\angle_{t}(V_{N},K_{t})\leq\frac{N\sqrt{5\pi(n-N)\log(n/\delta)}}{\sqrt{2}\delta(t-1)!}\Bigl{(}\frac{2}{\mathsf{relgap}}\Bigr{)}^{t-1}.$
Second, in Section 3.1.2, we prove
(10) $\tan\angle_{t}(V_{N},K_{\ell})\leq\frac{\tan\angle_{t}(V_{N},K_{t})}{\mathsf{Cheb}_{\ell-t}\big{(}1+2\frac{\lambda_{N+1}-\lambda_{N}}{\lambda_{n}-\lambda_{N+1}}\big{)}}.$
Plugging Eq. 9 into Eq. 10 then proves the inequality Eq. 6 of Theorem 3.1.
3.1.1 The case $\ell=t$. Proof of Eq. 9
Following [19, Theorem 3], the ability of a (single-vector) Krylov subspace to capture a subspace is closely related to properties of Vandermonde matrices.
Lemma 3.3.
For $\lambda_{1}<\dotsb<\lambda_{N}\leq\lambda_{N+1}\leq\dotsb\leq\lambda_{n}$ and $1\leq t\leq N$, consider the Vandermonde matrices
(11) $\mathsf{Van}_{N}=\begin{bmatrix}1&\lambda_{1}&\ldots&\lambda_{1}^{t-1}\\ 1&\lambda_{2}&\ldots&\lambda_{2}^{t-1}\\ \vdots&\vdots&\ddots&\vdots\\ 1&\lambda_{N}&\ldots&\lambda_{N}^{t-1}\end{bmatrix},\quad\mathsf{Van}_{\perp}=\begin{bmatrix}1&\lambda_{N+1}&\ldots&\lambda_{N+1}^{t-1}\\ 1&\lambda_{N+2}&\ldots&\lambda_{N+2}^{t-1}\\ \vdots&\vdots&\ddots&\vdots\\ 1&\lambda_{n}&\ldots&\lambda_{n}^{t-1}\end{bmatrix}.$
Then
$\lVert\mathsf{Van}_{\perp}\mathsf{Van}_{N}^{\dagger}\rVert\leq\frac{\sqrt{n-N}}{(t-1)!}\Bigl{(}\frac{2}{\mathsf{relgap}}\Bigr{)}^{t-1}$
with $\mathsf{relgap}$ defined as in Eq. 7.
Proof 3.4.
For an arbitrary vector $x=[x_{1},\dotsc,x_{t}]^{\mathsf{T}}\in\mathbb{R}^{t}$, we have
$\lVert\mathsf{Van}_{N}x\rVert^{2}=\sum_{i=1}^{N}p^{2}(\lambda_{i})\quad\text{and}\quad\lVert\mathsf{Van}_{\perp}x\rVert^{2}=\sum_{j=N+1}^{n}p^{2}(\lambda_{j}),$
with the polynomial $p(\lambda)=\sum_{k=1}^{t}x_{k}\lambda^{k-1}$. By the assumptions, $\mathsf{Van}_{N}$ has full column rank and
(12) $\lVert\mathsf{Van}_{\perp}\mathsf{Van}_{N}^{\dagger}\rVert^{2}=\max_{x\in\mathbb{R}^{t}}\frac{\lVert\mathsf{Van}_{\perp}x\rVert^{2}}{\lVert\mathsf{Van}_{N}x\rVert^{2}}=\max_{p\in\mathcal{P}_{t-1}}\frac{\sum_{j=N+1}^{n}p^{2}(\lambda_{j})}{\sum_{i=1}^{N}p^{2}(\lambda_{i})},$
where $\mathcal{P}_{t-1}$ denotes the vector space containing all polynomials of degree at most $t-1$. Choosing the Lagrange basis
$p_{i}(\lambda)=\prod_{\begin{subarray}{c}1\leq k\leq t\\ k\neq i\end{subarray}}\frac{\lambda-\lambda_{k}}{\lambda_{i}-\lambda_{k}}\in\mathcal{P}_{t-1},\quad i=1,\dotsc,t,$
we can express an arbitrary polynomial $p\in\mathcal{P}_{t-1}$ as $p(\lambda)=p(\lambda_{1})p_{1}(\lambda)+\dotsb+p(\lambda_{t})p_{t}(\lambda).$ By the Cauchy–Schwarz inequality,
(13) $p^{2}(\lambda)\leq\sum_{i=1}^{t}p^{2}(\lambda_{i})\sum_{i=1}^{t}p_{i}^{2}(\lambda)\leq\sum_{i=1}^{N}p^{2}(\lambda_{i})\Bigl{(}\sum_{i=1}^{t}\lvert p_{i}(\lambda)\rvert\Bigr{)}^{2}.$
To process the last term, we assume that $\lambda_{N+1}\leq\lambda\leq\lambda_{n}$, which implies $\lvert\lambda_{i}-\lambda_{k}\rvert/(\lambda-\lambda_{k})\geq\lvert i-k\rvert\cdot\mathsf{relgap}$ for $1\leq i\neq k\leq t$. In turn,
$\sum_{i=1}^{t}\lvert p_{i}(\lambda)\rvert=\sum_{i=1}^{t}\prod_{\begin{subarray}{c}1\leq k\leq t\\ k\neq i\end{subarray}}\frac{\lambda-\lambda_{k}}{\lvert\lambda_{i}-\lambda_{k}\rvert}\leq\frac{1}{\mathsf{relgap}^{t-1}}\sum_{i=1}^{t}\prod_{\begin{subarray}{c}1\leq k\leq t\\ k\neq i\end{subarray}}\frac{1}{\lvert i-k\rvert}=\Bigl{(}\frac{2}{\mathsf{relgap}}\Bigr{)}^{t-1}\frac{1}{(t-1)!},$
where the last equality utilizes the combinatorial identity
$\sum_{i=1}^{t}\prod_{\begin{subarray}{c}1\leq k\leq t\\ k\neq i\end{subarray}}\frac{1}{\lvert i-k\rvert}=\sum_{i=1}^{t}\frac{1}{(i-1)!(t-i)!}=\frac{1}{(t-1)!}\sum_{i=1}^{t}\binom{t-1}{i-1}=\frac{2^{t-1}}{(t-1)!}.$
Plugging this bound into (13) for $\lambda=\lambda_{j}$ gives
$\frac{\sum_{j=N+1}^{n}p^{2}(\lambda_{j})}{\sum_{i=1}^{N}p^{2}(\lambda_{i})}\leq\frac{n-N}{\bigl{(}(t-1)!\bigr{)}^{2}}\Bigl{(}\frac{2}{\mathsf{relgap}}\Bigr{)}^{2(t-1)}.$
Combined with Eq. 12, this completes the proof.
By leveraging the randomness of $\omega$, Proposition 2.11 allows us to prove Eq. 9.
Proof 3.5 (Proof of Eq. 9).
Because $V_{N}^{\mathsf{T}}\omega$ is a Gaussian random vector, the matrix $\mathrm{diag}(V_{N}^{\mathsf{T}}\omega)$ is invertible (almost surely). Using $V_{N}^{\mathsf{T}}K_{t}=\mathrm{diag}(V_{N}^{\mathsf{T}}\omega)\mathsf{Van}_{N}$, the invertibility of this matrix allows us to conclude that the Krylov basis $K_{t}=[\omega,B\omega,\dotsc,B^{t-1}\omega]$ has full rank:
(14) $\mathrm{rank}(K_{t})\geq\mathrm{rank}(V_{N}^{\mathsf{T}}K_{t})=\mathrm{rank}(\mathsf{Van}_{N})=t,$
where the last equality utilizes that the $N$ smallest eigenvalues of $B$ are mutually different. Letting $V_{\perp}$ denote an orthonormal basis of the invariant subspace associated with $\lambda_{N+1},\dotsc,\lambda_{n}$, we now apply Proposition 2.11 to estimate
(15) $\tan^{2}\angle_{t}(V_{N},K_{t})$ $=\lVert V_{\perp}^{\mathsf{T}}K_{t}(V_{N}^{\mathsf{T}}K_{t})^{\dagger}\rVert^{2}$
$=\lVert\mathrm{diag}(V_{\perp}^{\mathsf{T}}\omega)\mathsf{Van}_{\perp}(\mathrm{diag}(V_{N}^{\mathsf{T}}\omega)\mathsf{Van}_{N})^{\dagger}\rVert^{2}$
$\leq\frac{\sigma_{\max}^{2}(\mathrm{diag}(V_{\perp}^{\mathsf{T}}\omega))}{\sigma_{\min}^{2}(\mathrm{diag}(V_{N}^{\mathsf{T}}\omega))}\lVert\mathsf{Van}_{\perp}\mathsf{Van}_{N}^{\dagger}\rVert^{2}.$
Because $V_{\perp}^{\mathsf{T}}\omega$ and $V_{N}^{\mathsf{T}}\omega$ are independent Gaussian random vectors, we can use concentration and anti-concentration results for chi-squared random variables from [15, Lemma 1] and [19, Lemma 4], respectively, to conclude that
$\frac{\sigma_{\max}^{2}(\mathrm{diag}(V_{\perp}^{\mathsf{T}}\omega))}{\sigma_{\min}^{2}(\mathrm{diag}(V_{N}^{\mathsf{T}}\omega))}\leq\frac{5\pi N^{2}\log\bigl{(}n/\delta\bigr{)}}{2\delta^{2}}$
holds with probability at least $1-2\delta$. Combined with the bound on $\lVert\mathsf{Van}_{\perp}\mathsf{Van}_{N}^{\dagger}\rVert$ from Lemma 3.3, this completes the proof of Eq. 9.
3.1.2 The case $\ell>t$. Proof of Eq. 10
Our proof crucially relies on a simple but important observation from [19, (1.3)], which establishes the following equivalence between the single-vector Krylov subspace $\mathrm{span}\{K_{\ell}\}$ and a block Krylov subspace generated from a specific starting matrix:
(16) $\mathrm{span}\{K_{\ell}\}=\mathrm{span}\{\omega,B\omega,\dotsc,B^{\ell-1}\omega\}=\mathrm{span}\{K_{t},BK_{t},\dotsc,B^{\ell-t}K_{t}\},$
with $t\leq\ell\leq n$. The following lemma is a straightforward extension of known convergence results for Krylov subspaces to block Krylov subspaces.
Lemma 3.6.
For a symmetric matrix $B\in\mathbb{R}^{n\times n}$ with eigenvalues $\lambda_{1}<\lambda_{2}<\dotsb<\lambda_{N}<\lambda_{N+1}\leq\dotsb\leq\lambda_{n}$, let $V_{N}\in\mathbb{R}^{n\times N}$ be an orthonormal basis of the invariant subspace associated with $\lambda_{1},\dotsc,\lambda_{N}$. Choose integers $t,\ell$ such that $1\leq t\leq N$ and $t\leq\ell\leq n$. For an $n\times t$ matrix $\Omega$ such that $\mathrm{rank}(V_{N}^{\mathsf{T}}\Omega)=t$, let $K=[\Omega,B\Omega,\dotsc,B^{\ell-t}\Omega]\in\mathbb{R}^{n\times(\ell-t+1)t}$. Then
(17) $\tan\angle_{t}(V_{N},K)\leq\frac{\tan\angle_{t}(V_{N},\Omega)}{\mathsf{Cheb}_{\ell-t}\Bigl{(}1+2\frac{\lambda_{N+1}-\lambda_{N}}{\lambda_{n}-\lambda_{N+1}}\Bigr{)}},$
where $\mathsf{Cheb}_{\ell-t}$ is the Chebyshev polynomial Eq. 5 of degree $\ell-t$.
Proof 3.7.
Let $p(\lambda)=\alpha_{0}+\alpha_{1}\lambda+\dotsb\alpha_{\ell-t}\lambda^{\ell-t}$ be such that $p(\lambda)\not=0$ for $\lambda_{1}\leq\lambda\leq\lambda_{N}$. Defining the matrix $H_{K}=[\alpha_{0}I_{t},\alpha_{1}I_{t},\dotsc,\alpha_{\ell-t}I_{t}]^{\mathsf{T}}$, we have that
$V_{N}^{\mathsf{T}}KH_{K}=\sum_{i=0}^{\ell-t}\alpha_{i}V_{N}^{\mathsf{T}}B^{i}\Omega=\sum_{i=0}^{\ell-t}\alpha_{i}\Lambda_{N}^{i}V_{N}^{\mathsf{T}}\Omega=p(\Lambda_{N})V_{N}^{\mathsf{T}}\Omega,$
and an analogous formula holds when $V_{N}$ is replaced by $V_{\perp}$. By assumption, $p(\Lambda_{N})$ is invertible and thus $V_{N}^{\mathsf{T}}KH_{K}$ has rank $t$. Additionally, using that $\mathrm{span}\{KH_{K}\}\subset\mathrm{span}\{K\}$ and the expression for $\tan\angle_{t}$ from Proposition 2.11, we obtain that
$\tan\angle_{t}(V_{N},K)$ $\leq\tan\angle_{t}(V_{N},KH_{K})=\lVert V_{\perp}^{\mathsf{T}}KH_{K}(V_{N}^{\mathsf{T}}KH_{K})^{\dagger}\rVert$
$=\lVert p(\Lambda_{\perp})V_{\perp}^{\mathsf{T}}\Omega(V_{N}^{\mathsf{T}}\Omega)^{\dagger}p(\Lambda_{N})^{-1}\rVert$
$\leq\frac{\max\limits_{N+1\leq j\leq n}\lvert p(\lambda_{j})\rvert}{\min\limits_{1\leq i\leq N}\lvert p(\lambda_{i})\rvert}\lVert V_{\perp}^{\mathsf{T}}\Omega(V_{N}^{\mathsf{T}}\Omega)^{\dagger}\rVert=\frac{\max\limits_{N+1\leq j\leq n}\lvert p(\lambda_{j})\rvert}{\min\limits_{1\leq i\leq N}\lvert p(\lambda_{i})\rvert}\tan\angle_{t}(V_{N},\Omega).$
The proof is completed by setting $p(\lambda)=\mathsf{Cheb}_{\ell-t}\bigl{(}1+\frac{2(\lambda_{N+1}-\lambda)}{\lambda_{n}-\lambda_{N+1}}\bigr{)}$ and using that a Chebyshev polynomial is bounded by one in magnitude inside $[-1,1]$ and monotone outside that interval.
3.2 Convergence of Algorithm 1
In this subsection, we investigate the quality of the approximate null space produced by Algorithm 1 via applying the Rayleigh–Ritz procedure to the Krylov subspace. Unlike existing results on the block Lanczos method [16], we also present the convergence of lower-dimensional subspaces by bounding the principal angles $\angle_{t}$ for all $1\leq t\leq N$. We start with a result that establishes the quasi-optimality of Rayleigh–Ritz.
Lemma 3.8.
Under the setting of Theorem 2.6, let $V$ be the approximate null space basis returned after $\ell\geq N$ iterations of Algorithm 1. Let $\tilde{\lambda}_{t}^{(0)}$ be the $t$th smallest eigenvalue of $Z_{\ell}^{\mathsf{T}}A^{\mathsf{T}}AZ_{\ell}$, where $1\leq t\leq\ell$ and $Z_{\ell}$ is an orthonormal basis of $\mathrm{span}\{\omega,B\omega,\dotsc,B^{\ell-1}\omega\}$. Then for all $1\leq t\leq N$, there exists a matrix $W_{t}\in\mathbb{R}^{N\times t}$ with orthonormal columns such that
(18) $\lVert AVW_{t}\rVert^{2}\leq\tilde{\lambda}_{t}^{(0)}+\epsilon.$
Furthermore, if $t=N$ and $0<\epsilon<\underline{\sigma}^{2}-\tilde{\lambda}_{N}^{(0)}$, then
(19) $\lVert AV\rVert^{2}\leq\tilde{\lambda}_{N}^{(0)}+\frac{\epsilon^{2}}{\underline{\sigma}^{2}-\tilde{\lambda}_{N}^{(0)}-\epsilon}.$
Proof 3.9.
Let $V_{Z}$ and $V_{Z}W_{t}$ be two matrices with orthonormal columns, where their columns form orthonormal bases for the invariant subspaces associated with the $N$ and $t$ smallest eigenvalues of $Z_{\ell}^{\mathsf{T}}BZ_{\ell}$, respectively. By definition, $V=Z_{\ell}V_{Z}$. The gap-independent perturbation result in Proposition 2.4 allows us to conclude (18):
$\lVert AVW_{t}\rVert^{2}=\lVert(V_{Z}W_{t})^{\mathsf{T}}(Z_{\ell}^{\mathsf{T}}A^{\mathsf{T}}AZ_{\ell})(V_{Z}W_{t})\rVert\leq\tilde{\lambda}_{t}^{(0)}+\lVert Z_{\ell}^{\mathsf{T}}(B-A^{\mathsf{T}}A)Z_{\ell}\rVert\leq\tilde{\lambda}_{t}^{(0)}+\epsilon,$
where we used that $\tilde{\lambda}_{t}^{(0)}$ is the $t$th smallest eigenvalue of $Z_{\ell}^{\mathsf{T}}A^{\mathsf{T}}AZ_{\ell}$.
For Eq. 19, the result when $\ell=N$ is straightforward since $\lVert AV\rVert^{2}=\tilde{\lambda}_{N}^{(0)}$. To show the case $\ell>N$, we first note that eigenvalue interlacing implies $\tilde{\lambda}_{N+1}^{(0)}\geq\underline{\sigma}^{2}$. Because of $\|Z_{\ell}^{\mathsf{T}}BZ_{\ell}-Z_{\ell}^{\mathsf{T}}A^{\mathsf{T}}AZ_{\ell}\|\leq\epsilon<\underline{\sigma}^{2}-\tilde{\lambda}_{N}^{(0)}\leq\tilde{\lambda}_{N+1}^{(0)}-\tilde{\lambda}_{N}^{(0)}$, we can apply the second, gap-dependent result in Proposition 2.4 to establish (19).
By combining the eigenvalue repulsion result from Theorem 2.2, the Krylov subspace convergence result from Theorem 3.1 and the quasi-optimality of Rayleigh–Ritz from Lemma 3.8, we can conclude our main convergence result on the convergence of Algorithm 1.
Theorem 3.10.
Let $\overline{\sigma}$ and $\underline{\sigma}$ denote the largest and smallest nonzero singular values of a matrix $A\in\mathbb{R}^{m\times n}$ having null space dimension $N$. Let $V$ be the approximate null space basis returned by Algorithm 1 with $30<\epsilon<\underline{\sigma}^{2}/$ 3 and $N\leq\ell\leq n$. For every $1\leq t\leq N$, there exists a matrix $W_{t}\in\mathbb{R}^{N\times t}$ with orthonormal columns such that
(20) $\lVert AVW_{t}\rVert^{2}\leq\Bigl{(}\frac{\epsilon}{(\underline{\sigma}^{2}-\epsilon)^{1/2}}+\overline{\sigma}\tan\angle_{t}(V_{N},Z_{\ell})\Bigr{)}^{2}+\epsilon,$
holds with probability at least $1-3\delta$, where $\tan\angle_{t}(V_{N},Z_{\ell})$ is bounded according to Eq. 8. For $t=N$, when
(21) $\tan\angle_{N}(V_{N},Z_{\ell})\leq\frac{2\underline{\sigma}^{2}-5\epsilon}{3\overline{\sigma}(\underline{\sigma}^{2}-\epsilon)^{1/2}},$
the bound can be improved to
(22) $\lVert AV\rVert^{2}\leq\Bigl{(}\frac{\epsilon}{(\underline{\sigma}^{2}-\epsilon)^{1/2}}+\overline{\sigma}\tan\angle_{N}(V_{N},Z_{\ell})\Bigr{)}^{2}+\frac{9\epsilon^{2}}{5(\underline{\sigma}^{2}-\epsilon)}.$
Proof 3.11.
Using Lemma 3.8, it suffices to bound $\tilde{\lambda}_{t}^{(0)}$, the $t$th smallest eigenvalue of $Z_{\ell}^{\mathsf{T}}A^{\mathsf{T}}AZ_{\ell}$. Using Proposition 2.9, Theorem 2.6, and $\lvert\sin\angle\rvert\leq\lvert\tan\angle\rvert$, we derive
$\sqrt{\tilde{\lambda}_{t}^{(0)}}\leq\lVert AV_{N}\rVert+\lVert A\rVert\sin\angle_{t}(V_{N},Z_{\ell})\leq\frac{\epsilon}{(\underline{\sigma}^{2}-\epsilon)^{1/2}}+\overline{\sigma}\tan\angle_{t}(V_{N},Z_{\ell}).$
Combined with Eq. 18, this gives Eq. 20.
In order to prove Eq. 22, we first note that the condition Eq. 21 implies
$\underline{\sigma}^{2}-\epsilon-\tilde{\lambda}_{N}^{(0)}\geq\underline{\sigma}^{2}-\epsilon-\Bigl{(}\frac{\epsilon}{(\underline{\sigma}^{2}-\epsilon)^{1/2}}+\overline{\sigma}\tan\angle_{N}(V_{N},Z_{\ell})\Bigr{)}^{2}\geq\frac{5(\underline{\sigma}^{2}-\epsilon)}{9}>0.$
Thus, the condition for the bound Eq. 19 in Lemma 3.8 is met, which implies the desired result Eq. 22.
Remark 3.12.
Let $\kappa=\overline{\sigma}/\underline{\sigma}$ be the effective condition number of $A$, and $\epsilon=c\underline{\sigma}^{2}$, where $30<c<1/$ 3. Theorem 3.10 states that, with probability at least $1-3\delta$,
$\ell=\mathcal{O}\biggl{(}1+\frac{\log\bigl{(}(n\kappa)/(t\delta\epsilon)\bigr{)}}{\log(1+2\kappa^{-1})}\biggr{)}t$
iterations of Algorithm 1 are needed to partially approximate the null space of $A$ in the sense of
$\tan\angle_{t}(V_{N},Z_{\ell})\leq\frac{\epsilon}{\sqrt{6}\overline{\sigma}\underline{\sigma}}\leq\frac{2\underline{\sigma}^{2}-5\epsilon}{3\overline{\sigma}(\underline{\sigma}^{2}-\epsilon)^{1/2}},\quad\text{and}\quad\lVert AVW_{t}\rVert^{2}\leq\frac{17\epsilon}{9},$
where we used Stirling’s formula and a known bound for Chebyshev polynomials in [28, Section 4.4.1]. Let $BZ_{\ell}=Z_{\ell}T_{\ell}+Q_{\ell+1}E_{\ell+1}$ represent the Lanczos decomposition computed by Algorithm 1. Then, $T_{\ell}-3\epsilon I$ has at least $t$ negative eigenvalues, all originating from Ritz vectors close to the null space. This assertion stems from the fact that the $(N+1)$st smallest eigenvalue of $B$ is at least $\underline{\sigma}^{2}=\epsilon/c>3\epsilon$. Moreover, for $t=N$, the error bound improves to
$\lVert AV\rVert^{2}\leq\Bigl{(}\frac{\epsilon}{(\underline{\sigma}^{2}-\epsilon)^{1/2}}+\frac{\epsilon}{\sqrt{6}\underline{\sigma}}\Bigr{)}^{2}+\frac{9\epsilon^{2}}{5(\underline{\sigma}^{2}-\epsilon)}\leq\frac{161\epsilon^{2}}{30\underline{\sigma}^{2}}.$
Our theoretical results highlight a crucial distinction between the convergence behaviors of the single-vector and block Lanczos methods. In contrast to the block Lanczos method [16], where Ritz vectors converge collectively, the single-vector Lanczos method exhibits individual convergence of Ritz vectors. This individual convergence allows for deflation and adaptive detection of the null space dimension, which is particularly advantageous in the typical scenario where prior information about $N$ is unavailable.
4 Practical techniques for implementation
In this section, we discuss several improvements and implementation details to make Algorithm 1 competitive.
4.1 Restarting
Algorithm 1 needs to store the orthonormal basis $Z_{\ell}$ of the Krylov subspace to allow for reorthogonalization (see below) and Ritz vector computations. As the number of iterations $\ell$ required for convergence increases, storing $Z_{\ell}$ becomes challenging due to memory limitations. Additionally, the cost of Rayleigh–Ritz extraction grows significantly with $\ell$. To address these issues, we employ a restart mechanism, akin to the Krylov–Schur method [32, 37].
To conveniently describe our restarting mechanism, we use the superscripts $\cdot^{(\mathrm{old})}$ and $\cdot^{(\mathrm{new})}$ to denote variables before and after restarting, respectively. We assume that the dimension of the Krylov subspace before restarting is $\ell d$ and that $\ell_{0}d$ Ritz pairs are retained after restarting, where $\ell_{0}<\ell$ and $d$ represents the block size of the initial matrix $\Omega$.
Suppose we have a Lanczos decomposition of the form
(23) $BZ_{\ell}^{(\mathrm{old})}=Z_{\ell}^{(\mathrm{old})}T_{\ell}^{(\mathrm{old})}+Q_{\ell+1}^{(\mathrm{old})}E_{\ell+1}^{(\mathrm{old})},$
where $[Z_{\ell}^{(\mathrm{old})},Q_{\ell+1}^{(\mathrm{old})}]$ is a matrix with orthonormal columns, $E_{\ell+1}^{(\mathrm{old})}=[0,\Gamma_{\ell}^{(\mathrm{old})}]$, and
(24) $T_{\ell}^{(\mathrm{old})}=\begin{bmatrix}\Theta^{(\mathrm{old})}&F^{(\mathrm{old})}&&&&\\ (F^{(\mathrm{old})})^{\mathsf{T}}&\Xi_{\ell_{0}+1}^{(\mathrm{old})}&\Gamma_{\ell_{0}+1}^{(\mathrm{old})}&&&\\ &(\Gamma_{\ell_{0}+1}^{(\mathrm{old})})^{\mathsf{T}}&\Xi_{\ell_{0}+2}^{(\mathrm{old})}&\Gamma_{\ell_{0}+2}^{(\mathrm{old})}&&\\ &&\ddots&\ddots&\ddots&\\ &&&(\Gamma_{\ell-2}^{(\mathrm{old})})^{\mathsf{T}}&\Xi_{\ell-1}^{(\mathrm{old})}&\Gamma_{\ell-1}^{(\mathrm{old})}\\ &&&&(\Gamma_{\ell-1}^{(\mathrm{old})})^{\mathsf{T}}&\Xi_{\ell}^{(\mathrm{old})}\\ \end{bmatrix}$
is a block tridiagonal matrix, partitioned such that $\Theta^{(\mathrm{old})}\in\mathbb{R}^{\ell_{0}d\times\ell_{0}d}$ contains the information from the previous cycle. To restart and keep $\ell_{0}d$ Ritz vectors, we apply the Rayleigh–Ritz procedure:
$T_{\ell}^{(\mathrm{old})}V_{Z}^{(\mathrm{new})}=V_{Z}^{(\mathrm{new})}\Theta^{(\mathrm{new})}\quad\text{and}\quad Z_{\ell_{0}}^{(\mathrm{new})}=Z_{\ell}^{(\mathrm{old})}V_{Z}^{(\mathrm{new})}\in\mathbb{R}^{n\times\ell_{0}d},$
where $V_{Z}^{(\mathrm{new})}\in\mathbb{R}^{\ell d\times\ell_{0}d}$ is an orthonormal basis of the invariant subspace corresponding to the $\ell_{0}d$ smallest eigenvalues of $T_{\ell}^{(\mathrm{old})}$. Multiplying $V_{Z}^{(\mathrm{new})}$ from the right to the Lanczos decomposition Eq. 23 leads to
$BZ_{\ell_{0}}^{(\mathrm{new})}=Z_{\ell_{0}}^{(\mathrm{new})}\Theta^{(\mathrm{new})}+Q_{\ell+1}^{(\mathrm{old})}E_{\ell+1}^{(\mathrm{old})}V_{Z}^{(\mathrm{new})}.$
We restart the block Lanczos method by defining the quantities
$Q_{\ell_{0}+1}^{(\mathrm{new})}=Q_{\ell+1}^{(\mathrm{old})}\quad\text{and}\quad Z_{\ell_{0}+1}^{(\mathrm{new})}=[Z_{\ell_{0}}^{(\mathrm{new})},Q_{\ell_{0}+1}^{(\mathrm{new})}]\in\mathbb{R}^{n\times(\ell_{0}+1)d},$
$F^{(\mathrm{new})}=(E_{\ell+1}^{(\mathrm{old})}V_{Z}^{(\mathrm{new})})^{\mathsf{T}}\quad\text{and}\quad\Xi_{\ell_{0}+1}^{(\mathrm{new})}=(Q_{\ell_{0}+1}^{(\mathrm{new})})^{\mathsf{T}}BQ_{\ell_{0}+1}^{(\mathrm{new})},$
$\bigl{(}I-Z_{\ell_{0}+1}^{(\mathrm{new})}(Z_{\ell_{0}+1}^{(\mathrm{new})})^{\mathsf{T}}\bigr{)}BQ_{\ell_{0}+1}^{(\mathrm{new})}=Q_{\ell_{0}+2}^{(\mathrm{new})}\Gamma_{\ell_{0}+1}^{(\mathrm{new})},$
where the last line denotes a QR decomposition. After restarting, the new Lanczos decomposition takes the form
$BZ_{\ell_{0}+1}^{(\mathrm{new})}$ $=[BZ_{\ell_{0}}^{(\mathrm{new})},BQ_{\ell_{0}+1}^{(\mathrm{new})}]$
$=[Z_{\ell_{0}}^{(\mathrm{new})},Q_{\ell_{0}+1}^{(\mathrm{new})}]\begin{bmatrix}\Theta^{(\mathrm{new})}&(E_{\ell+1}^{(\mathrm{old})}V_{Z}^{(\mathrm{new})})^{\mathsf{T}}\\ E_{\ell+1}^{(\mathrm{old})}V_{Z}^{(\mathrm{new})}&\Xi_{\ell_{0}+1}^{(\mathrm{new})}\end{bmatrix}+Q_{\ell_{0}+2}^{(\mathrm{new})}[O,\Gamma_{\ell_{0}+1}^{(\mathrm{new})}]$
$=Z_{\ell_{0}+1}^{(\mathrm{new})}\begin{bmatrix}\Theta^{(\mathrm{new})}&F^{(\mathrm{new})}\\ (F^{(\mathrm{new})})^{\mathsf{T}}&\Xi_{\ell_{0}+1}^{(\mathrm{new})}\end{bmatrix}+Q_{\ell_{0}+2}^{(\mathrm{new})}E_{\ell_{0}+2}^{(\mathrm{new})},$
where $E_{\ell_{0}+2}^{(\mathrm{new})}=[O,\Gamma_{\ell_{0}+1}^{(\mathrm{new})}]$, and the relationship
$E_{\ell+1}^{(\mathrm{old})}V_{Z}^{(\mathrm{new})}=(Q_{\ell+1}^{(\mathrm{old})})^{\mathsf{T}}(BZ_{\ell_{0}}^{(\mathrm{new})}-Z_{\ell_{0}}^{(\mathrm{new})}\Theta^{(\mathrm{new})})=(Q_{\ell_{0}+1}^{(\mathrm{new})})^{\mathsf{T}}BZ_{\ell_{0}}^{(\mathrm{new})}$
is used.
4.2 Partial reorthogonalization
In floating-point arithmetic, round-off error compromises the orthogonality of the basis computed by the block Lanczos method, leading to the emergence of multiple converged Ritz values, so called ghost eigenvalues. To obtain reliable information about the null space and its dimension, it is imperative to maintain the orthogonality of Lanczos vectors with reorthogonalization strategies.
Commonly employed reorthogonalization schemes include selective reorthogonalization [24], partial reorthogonalization [31], and full reorthogonalization. In our implementation, we use partial reorthogonalization, which is cheaper than full reorthogonalization and more effective than selective reorthogonalization [30]. Following the strategies described in [33, 12], we define a $W$-recurrence to monitor loss of orthogonality. For brevity, we omit the superscripts $\cdot^{(\mathrm{old})}$ and $\cdot^{(\mathrm{new})}$ regarding restarting. Additionally, we focus on the reorthogonalization of the new basis vector $Q_{\ell+1}$ versus $Z_{\ell}=[Q_{1},\dotsc,Q_{\ell}]$.
Supposing that $\ell_{0}d$ Ritz vectors were retained in the last restart, and
$\Theta=\mathrm{diag}\{\Theta_{1},\dotsc,\Theta_{\ell_{0}}\}\quad\text{and}\quad F=[F_{1}^{\mathsf{T}},\dotsc,F_{\ell_{0}}^{\mathsf{T}}]^{\mathsf{T}},$
the recurrence Eq. 23 in finite precision can be written as
(25) $BQ_{i}$ $=Q_{i}\Theta_{i}+Q_{\ell_{0}}F_{i}^{\mathsf{T}}+\Delta_{i}\quad\text{for}\quad 1\leq i\leq\ell_{0},$
$BQ_{\ell_{0}+1}$ $=\sum_{i=1}^{\ell_{0}}Q_{i}F_{i}+Q_{\ell_{0}+1}\Xi_{\ell_{0}+1}+Q_{\ell_{0}+2}\Gamma_{\ell_{0}+1}^{\mathsf{T}}+\Delta_{\ell_{0}+1},$
$BQ_{j}$ $=Q_{j-1}\Gamma_{j-1}+Q_{j}\Xi_{j}+Q_{j+1}\Gamma_{j}^{\mathsf{T}}+\Delta_{j}\quad\text{for}\quad\ell_{0}+2\leq j\leq\ell,$
where the matrices denoted with $\Delta$ account for round-off errors. Following [12, 23], we may assume that the matrices $Q_{k}$ have orthonormal columns and
(26) $\lVert\Delta_{k}\rVert\leq d\sqrt{n}\mathbf{u}\lVert B\rVert=:\mathbf{u}_{d}$
for $11\leq k\leq\ell+$ 1, where $\mathbf{u}$ denotes the machine precision. Different from full orthogonalization, we allow $Q_{i}$ and $Q_{j}$ to be not (numerically) orthogonal to each other for $i\not=j$. The basic idea of partial reorthogonalization is to monitor this loss of orthogonality by examining $\lVert Q_{k}^{\mathsf{T}}Q_{\ell+1}\rVert$ for $1\leq k\leq\ell$. If this value exceeds a certain threshold $\mathbf{u}_{\mathsf{RO}}$, typically chosen as $\mathbf{u}_{\mathsf{RO}}=\sqrt{\mathbf{u}}$, we perform reorthogonalization. Because computing $\lVert Q_{k}^{\mathsf{T}}Q_{\ell+1}\rVert$ is computationally expensive, we instead aim to calculate and use an upper bound for this value. Reorthogonalization is not performed if the upper bound is below $\mathbf{u}_{\mathsf{RO}}$, potentially cutting down computational costs while maintaining a decent level of orthogonality.
To derive an upper bound for $\lVert Q_{k}^{\mathsf{T}}Q_{\ell+1}\rVert$, we take $j=\ell$ in the third recurrence in Eq. 25. Multiplying with $Q_{k}^{\mathsf{T}}$ from the left and rearranging terms yields the inequality
$\lVert Q_{k}^{\mathsf{T}}Q_{\ell+1}\rVert$ $\leq\lVert\Gamma_{\ell}^{-1}\rVert\bigl{\lVert}Q_{k}^{\mathsf{T}}BQ_{\ell}-Q_{k}^{\mathsf{T}}Q_{\ell-1}\Gamma_{\ell-1}-Q_{k}^{\mathsf{T}}Q_{\ell}\Xi_{\ell}-Q_{k}^{\mathsf{T}}\Delta_{\ell}\bigr{\rVert}$
$\leq\lVert\Gamma_{\ell}^{-1}\rVert\bigl{(}\lVert Q_{k}^{\mathsf{T}}BQ_{\ell}\rVert+\lVert Q_{k}^{\mathsf{T}}Q_{\ell-1}\rVert\lVert\Gamma_{\ell-1}\rVert+\lVert Q_{k}^{\mathsf{T}}Q_{\ell}\rVert\lVert\Xi_{\ell}\rVert+\lVert Q_{k}\rVert\lVert\Delta_{\ell}\rVert\bigr{)}.$
For $1\leq k\leq\ell_{0}$, the term $\lVert Q_{k}^{\mathsf{T}}BQ_{\ell}\rVert$ is bounded by using the symmetry of $B$ and the first recurrence in Eq. 25:
$\lVert Q_{k}^{\mathsf{T}}BQ_{\ell}\rVert$ $=\lVert Q_{\ell}^{\mathsf{T}}BQ_{k}\rVert\leq\lVert Q_{k}^{\mathsf{T}}Q_{\ell}\rVert\lVert\Theta_{k}\rVert+\lVert Q_{\ell_{0}}^{\mathsf{T}}Q_{\ell}\rVert\lVert F_{k}\rVert+\lVert Q_{\ell}\rVert\lVert\Delta_{k}\rVert.$
Combining these two inequalities and using Eq. 26 yields
$\lVert Q_{k}^{\mathsf{T}}Q_{\ell+1}\rVert\leq\lVert\Gamma_{\ell}^{-1}\rVert\bigl{(}\lVert Q_{k}^{\mathsf{T}}Q_{\ell}\rVert(\lVert\Theta_{k}\rVert+\lVert\Xi_{\ell}\rVert)+\lVert Q_{\ell_{0}}^{\mathsf{T}}Q_{\ell}\rVert\lVert F_{k}\rVert+\lVert Q_{k}^{\mathsf{T}}Q_{\ell-1}\rVert\lVert\Gamma_{\ell-1}\rVert+2\mathbf{u}_{d}\bigr{)}.$
We keep track of the recursive inequalities above using a strictly upper triangular matrix $W$, such that each entry $w_{i,j}$ is an upper bound on $\lVert Q_{i}^{\mathsf{T}}Q_{j}\rVert$ for $11\leq i<j\leq\ell+$ 1. Assuming that columns $1,\ldots,\ell$ have been determined, the entries $1,\ldots,\ell$ of column $\ell+1$ are determined by the recurrences above as follows:
$w_{k,\ell+1}=\left\{\begin{aligned} \lVert\Gamma_{\ell}^{-1}\rVert\Bigl{(}&w_{\ell_{0}+1,\ell}(\lVert\Theta_{k}\rVert+\lVert\Xi_{\ell}\rVert)+w_{\ell_{0},\ell}\lVert F_{k}\rVert\\ &+w_{k,\ell-1}\lVert\Gamma_{\ell-1}\rVert+2\mathbf{u}_{d}\Bigr{)},&\!\!\!\!\!\!\!\!\text{for }1\leq k\leq\ell_{0}.\\ \lVert\Gamma_{\ell}^{-1}\rVert\Bigl{(}&w_{k,\ell}(\lVert\Xi_{k}\rVert+\lVert\Xi_{\ell}\rVert)+\sum_{1\leq i\leq\ell_{0}}w_{i,\ell}\lVert F_{i}\rVert\\ &+w_{k+1,\ell}\lVert\Gamma_{k}\rVert+w_{k,\ell-1}\lVert\Gamma_{\ell-1}\rVert+2\mathbf{u}_{d}\Bigr{)},&\!\!\!\!\!\!\text{for }k=\ell_{0}+1.\\ \lVert\Gamma_{\ell}^{-1}\rVert\Bigl{(}&w_{k,\ell}(\lVert\Xi_{k}\rVert+\lVert\Xi_{\ell}\rVert)+w_{k+1,\ell}\lVert\Gamma_{k}\rVert\\ &+w_{k-1,\ell}\lVert\Gamma_{k-1}\rVert+w_{k,\ell-1}\lVert\Gamma_{\ell-1}\rVert+2\mathbf{u}_{d}\Bigr{)},&\text{for }\ell_{0}+2\leq k\leq\ell.\end{aligned}\right.$
Note that the computation of the column $\ell+1$ only involves the previous two columns $\ell-1$ and $\ell$, consistent with the three-term recurrence of the Lanczos method. The entries of the initial two columns $\ell_{0}+1$ and $\ell_{0}+2$ of $W$ are set to $\mathbf{u}_{d}$ because $Q_{\ell_{0}+1}$ and $Q_{\ell_{0}+2}$ are reorthogonalized versus $Z_{\ell_{0}}$ immediately after restarting. We monitor the loss of orthogonality using $W$. If $w_{k,\ell+1}>\mathbf{u}_{\mathsf{RO}}$ for some $k$, we reorthogonalize $Q_{\ell}$ and $Q_{\ell+1}$ versus $Z_{\ell-1}$ and set the strictly upper triangular entries of the corresponding columns $\ell$ and $\ell+1$ of $W$ to $\mathbf{u}_{d}$.
4.3 Pseudocode
Algorithm 2 incorporates the improvements to Algorithm 1 discussed in this section. For simplicity, we assume that $\dim\mathcal{K}$ is an integer multiple of $d$. The pseudocode uses Matlab’s colon notation to denote submatrices.
Input: Matrix $A\in\mathbb{R}^{m\times n}$ with $m\geq n$ and an unknown number of zero singular values. Perturbation parameter $\epsilon>0$. Block size $d$. Maximum allowed Krylov subspace dimension $\dim\mathcal{K}$.
Output: Orthonormal basis $V$ approximating the null space of $A$.
1 Set $B=A^{\mathsf{T}}A+\epsilon D$, with a diagonal matrix $D$ having diagonal entries uniformly i.i.d. in $[0,1]$;
2 Draw Gaussian random matrix $\Omega\in\mathbb{R}^{n\times d}$;
3 Set $\mathbf{u}_{d}=d\sqrt{n}\mathbf{u}\lVert B\rVert$ and $\ell_{0}=0$;
4 Compute $Q_{1}$ as an orthonormal basis of $\Omega$;
5 while not converged do
6       Compute $Q_{\ell_{0}+1}^{\prime}=BQ_{\ell_{0}+1}$ and $\Xi_{\ell_{0}+1}=Q_{\ell_{0}+1}^{\mathsf{T}}Q_{\ell_{0}+1}^{\prime}$;
7       Compute QR decomposition $(I-Z_{\ell_{0}+1}Z_{\ell_{0}+1}^{\mathsf{T}})Q_{\ell_{0}+1}^{\prime}=Q_{\ell_{0}+2}\Gamma_{\ell_{0}+1}$;
8       Set $Z_{\ell_{0}+2}=[Z_{\ell_{0}+1},Q_{\ell_{0}+2}]$;
9       for $\ell=\ell_{0}+2,\dotsc,\dim\mathcal{K}/d$ do
10             Compute $Q_{\ell}^{\prime}=BQ_{\ell}-Q_{\ell-1}\Gamma_{\ell-1}$ and $\Xi_{\ell}=Q_{\ell}^{\mathsf{T}}Q_{\ell}^{\prime}$;
11             Compute QR decomposition $Q_{\ell}^{\prime}-Q_{\ell}\Xi_{\ell}=Q_{\ell+1}\Gamma_{\ell}$;
12             Perform partial reorthogonalization as discussed in Section 4.2 and set $Z_{\ell}=[Z_{\ell-1},Q_{\ell}]$ ;
13
14       end for
15      Form block tridiagonal matrix $T_{\ell}$ defined in Eq. 24 and compute its eigenvalue decomposition $T_{\ell}V_{Z}=V_{Z}\Theta$ with eigenvalues in ascending order;
16       Let $N_{\ell}$ be the number of eigenvalues of $T_{\ell}$ below $3\epsilon$ and set $\ell_{0}=(N_{\ell}+\dim\mathcal{K})/2$;
17       Set $\Theta=\Theta(1:\ell_{0},1:\ell_{0})$ and compute $Z_{\ell_{0}}=Z_{\ell}V_{Z}(:,1:\ell_{0})$;
18       Compute QR decomposition $(I-Z_{\ell_{0}}Z_{\ell_{0}}^{\mathsf{T}})Q_{\ell+1}=Q_{\ell_{0}+1}\Gamma_{\ell_{0}}$;
19       Set $F=([O,\Gamma_{\ell_{0}}]V_{Z})^{\mathsf{T}}$ and $Z_{\ell_{0}+1}=[Z_{\ell_{0}},Q_{\ell_{0}+1}]$;
20
21 end while
22return $V=Z_{\ell}V_{Z}(:,1:N_{\ell})$;
Algorithm 2 Improved randomized small-block Lanczos
4.4 Preconditioning
In Remark 3.12, we established that the number of iterations $\ell$ needed for convergence depends on the effective condition number $\kappa=\overline{\sigma}/\underline{\sigma}$. Preconditioning can be used to reduce this number.
4.4.1 Inner preconditioning
Given an $m\times m$ symmetric positive preconditioner $P_{L}$ for $AA^{\mathsf{T}}$, one may replace $B=A^{\mathsf{T}}A+\epsilon D$ by
$B_{L}=A^{\mathsf{T}}P_{L}^{-1}A+\epsilon D$
within the Lanczos method. Besides maintaining the null space (the null spaces of $A^{\mathsf{T}}P_{L}^{-1}A$ and $A$ are identical), the effective condition number is improved to $\kappa=\lVert(A^{\mathsf{T}}P_{L}^{-1}A)(A^{\mathsf{T}}P_{L}^{-1}A)^{\dagger}\rVert^{1/2}$. For $m\gg n$, using such a preconditioner may be expensive because $P_{L}$ is an $m\times m$ matrix.
4.4.2 Outer preconditioning
Given an invertible matrix $P_{R}\in\mathbb{R}^{n\times n}$ such that $P_{R}P_{R}^{\mathsf{T}}$ is a preconditioner for $A^{\mathsf{T}}A$, one may replace $B=A^{\mathsf{T}}A+\epsilon D$ by
$B_{R}=P_{R}^{-1}A^{\mathsf{T}}AP_{R}^{-\mathsf{T}}+\epsilon D$
within the Lanczos method. This transformation improves the effective condition number to $\kappa=\lVert(AP_{R}^{-\mathsf{T}})(AP_{R}^{-\mathsf{T}})^{\dagger}\rVert$. However, unlike inner preconditioning, this method necessitates the application of both $P_{R}^{-1}$ and $P_{R}^{-\mathsf{T}}$ in a single Lanczos step. Moreover, an additional step of $P_{R}^{-\mathsf{T}}V$ is necessary to post-process the output because $V$ approximates the null space of $AP_{R}^{-\mathsf{T}}$ instead of $A$.
5 Numerical experiments
In this section, we present some numerical results to support the theoretical results and demonstrate the efficiency of our method. All numerical experiments [^1] Scripts to reproduce numerical results are publicly available at https://github.com/nShao678/Nullspace-code in this section are implemented in Matlab 2022b and executed with an AMD Ryzen 9 6900HX Processor (8 cores, 3.3–4.9 GHz) and 32 GB of RAM.
5.1 Benchmark examples
For our experiments, we will consider two applications of null space computation: (1) Counting the connected components of an undirected graph and (2) cohomology computation.
For an undirected graph $G$, the null space $V_{G}$ of its graph Laplacian $L_{G}$ contains information about its connectivity. Specifically, the dimension of $V_{G}$ is the number of connected components, and the nodes in each connected component can be computed from a null space basis. As a specific example for a graph, we consider the mathematicians’ collaboration network from MathSciNet [26]. As this graph is connected, the null space of $L_{G}$ is trivial. To create a more interesting scenario, we consider the sub-graph $G_{0}$ obtained by first deleting nodes of degree less than $N_{0}$ for a fixed positive integer $N_{0}$ (along with the corresponding edges) and then deleting isolated nodes. Table 2 provides statistics in the resulting graphs for $N_{0}=3$ and $N_{0}=10$. As the graph Laplacian $A$ is symmetric positive semi-definite, we will replace $B=A^{\mathsf{T}}A+\epsilon D$ by $B=A+\epsilon D$ in line 1 of Algorithm 2 for these examples. Moreover, in order to accelerate convergence, we use outer preconditioning (see Section 4.4.2) with the preconditioner $P_{R}$ returned by the incomplete Cholesky decomposition of $A$, using the Matlab function
[…]
For $N_{0}=10$, this improves the effective condition number $\kappa=\overline{\sigma}/\underline{\sigma}$ from $6227.1$ to $23.2$.
Table 2: Statistics on graph Laplacians for MathSciNet collaboration network after preprocessing.
$N_{0}$nodes #connected components
3 168692 122
10 39890 66
For cohomology computations, we select the matrices from the SuiteSparse collection [5] listed in Table 3.
Table 3: SuiteSparse collection matrices related to cohomology computations.
Name #rows #columns nnz type nullity
Franz7 10164 1740 40424 Rings 113
Franz8 16728 7176 100368 Rings 1713
Franz9 19588 4164 97508 Rings 619
Franz10 19588 4164 97508 Rings 619
GL7d12 8899 1019 37519 $\mathrm{GL}_{7}(\mathbb{Z})$ 59
GL7d13 47271 8899 356232 $\mathrm{GL}_{7}(\mathbb{Z})$ 961
GL7d14 171375 47271 1831183 $\mathrm{GL}_{7}(\mathbb{Z})$ 7939
5.2 Single-vector Lanczos
We first assess the convergence of the single-vector Lanczos method with full reorthogonalization and without restarting for the graph Laplacian with $N_{0}=10$ from Section 5.1, utilizing the incomplete Cholesky preconditioner Eq. 27.
We varied the perturbation level $\epsilon$ from $4^{-4}$ to $4^{-10}$ and ran single-vector Lanczos method for $\ell=1000$ iterations to ensure convergence. Table 4 contains $\lambda_{N}(T_{\ell})$, the $N$th smallest eigenvalue of $T_{\ell}$, and norms $\lVert V^{\mathsf{T}}AV\rVert$ and $\lVert AV\rVert$. As expected, $\lambda_{N}(T_{\ell})$ and $\lVert V^{\mathsf{T}}AV\rVert$ are observed to be of the order $\mathcal{O}(\epsilon)$ and $\mathcal{O}(\epsilon^{2})$, respectively.
Table 4: Convergence of null space, where $\epsilon$ is taken as $4^{-k}$ for $k=4,\dotsc,10$.
$\epsilon$ 3.9e-03 9.8e-04 2.4e-04 6.1e-05 1.5e-05 3.8e-06 9.5e-07
$\lambda_{N}(T_{\ell})$ 3.7e-03 9.2e-04 2.3e-04 5.8e-05 1.4e-05 3.6e-06 9.0e-07
$\lVert AV\rVert$ 4.7e-03 1.2e-03 2.9e-04 7.4e-05 1.8e-05 4.6e-06 1.1e-06
$\lVert V^{\mathsf{T}}AV\rVert$ 1.5e-06 9.2e-08 5.7e-09 3.6e-10 2.2e-11 1.4e-12 8.7e-14
The result of Theorem 3.10 predicts that the Ritz values of the single-vector Lanczos method converge individually to the perturbed zero eigenvalues, that is, some converge faster than others.. The numerical results shown in Fig. 1 for $\epsilon=4^{-10}$ confirm this result, showing that the number of converged Ritz values grows steadily, a Ritz value converges approximately every $10$ steps of the Lanczos method.
Figure 1: Convergence history of single-vector Lanczos with $\epsilon=4^{-10}$ for graph Laplacian: Number of converged Ritz values vs. number of iterations.
5.3 Small-block Lanczos method with restarting and partial reorthogonalization
We now investigate the performance of Algorithm 2, the small-block Lanczos method with restarting and partial reorthogonalization. For all remaining experiments, we set the perturbation parameter to $\epsilon=10^{-3}$.
5.3.1 Comparison with null
First, we compare Algorithm 2 with the Matlab built-in function null, which is based on the SVD, for the first six matrices from Table 3. The obtained numerical results are presented in Table 5. For all parameter choices, the computed nullity matches the actual nullity from Table 3. It can be observed that restarting is very effective at limiting the dimension of the Krylov subspace. Although increasing the block size increases the number of matrix-vector multiplications, this increase is, for smaller $d$, outpaced by the higher efficiency of block matrix-vector multiplications. Once the block size becomes too large, both the iteration numbers and execution times start increasing. Based on our experiments, a suitable strategy for choosing the block size is $d=\dim\mathcal{K}/16$, where $\dim\mathcal{K}$ is the maximum dimension of the Krylov subspace before restarting. Algorithm 2 is faster than null for all examples, and the difference becomes more significant as the matrix size increases.
Table 5: Number of matrix vector multiplications and execution time of Algorithm 2 applied to cohomology computations. $\dim\mathcal{K}$ is the maximum dimension of Krylov subspace before restarting. The last column refers to Matlab’s built-in function null.
Block size $d=2$ $d=4$ $d=8$ $d=16$ $d=32$ null
##\##matvec 2520 2616 2824 3232 3904 $\times$
Time (s) 0.9023 0.7106 0.5412 0.6125 0.6515 1.5449
Block size $d=16$ $d=32$ $d=64$ $d=128$ $d=256$ null
##\##matvec 22848 22880 23808 25216 26880 $\times$
Time (s) 96.83 77.53 71.26 69.55 77.29 82.47
Block size $d=8$ $d=16$ $d=32$ $d=64$ $d=128$ null
##\##matvec 12496 13008 12960 14528 16000 $\times$
Time (s) 17.68 13.63 11.11 9.852 9.572 21.47
Block size $d=8$ $d=16$ $d=32$ $d=64$ $d=128$ null
##\##matvec 12440 12784 12960 14464 16000 $\times$
Time (s) 17.54 12.81 10.43 9.74 9.028 21.15
Block size $d=1$ $d=2$ $d=4$ $d=8$ $d=16$ null
##\##matvec 1439 1482 1504 1720 1920 $\times$
Time (s) 0.2188 0.1858 0.1545 0.1752 0.2387 0.6476
Block size $d=16$ $d=32$ $d=64$ $d=128$ $d=256$ null
##\##matvec 17552 18176 19072 19712 23040 $\times$
Time (s) 55.20 44.71 38.81 37.29 41.50 212.2
5.3.2 Comparison with large-block Lanczos
Traditionally, the block Lanczos method with restarting is used with a block size that is slightly larger than the nullity. Obviously, when memory is a constraint, the small-block or even single-vector Lanczos methods with restarting are preferred. Even when memory is not a constraint, using the small-block version can be beneficial. To illustrate this, we consider two examples: the matrix “Franz9” from cohomology computations and the graph Laplacian with $N_{0}=3$, which have nullities $619$ and $122$, respectively. We compare the performance of the small-block and large-block Lanczos methods. For the large-block Lanczos method, we set the maximum dimension of the Krylov subspace before restarting to be approximately four times the nullity, i.e., $2560$ and $512$, respectively.
As shown in Table 6, the large-block Lanczos method requires more matrix-vector multiplications and execution time compared to the small-block version. The nullity has been computed correctly by all methods. One advantage of the large-block Lanczos method is that it does not require small perturbations such as $B=A^{\mathsf{T}}A+\epsilon D$ or $B=A+\epsilon D$, thanks to the use of a sufficiently large-block size. While the absence of perturbations yields more accurate results, the overall computational cost remains unaffected because the additional computational cost from applying the diagonal perturbation matrix $D$ is negligible.
Table 6: Number of matrix vector multiplications, execution time, and null space residual of Algorithm 2 applied to “Franz9” and the graph Laplacian with $N_{0}=3$. $\dim\mathcal{K}$ is the maximum dimension of Krylov subspace before restarting. The last column refers to the block Lanczos method applied to the original matrix (without diagonal perturbations). The algorithm is stopped once $\lVert AV\rVert\leq 10^{-4}$ for “Franz9” or $\lVert AV\rVert\leq 1.2\times 10^{-3}$ for the graph Laplacian is detected before restarting.
Small-block Lanczos Block Lanczos
Block size $d=40$ $d=80$ $d=160$ $d=320$ $d=640$ $d=640$
##\##matvec 7960 6960 7200 8960 10240 10240
Time (s) 13.57 10.27 9.925 12.32 11.23 11.38
$\lVert AV\rVert$ 9.6e-05 9.6e-05 1.0e-04 9.6e-05 9.6e-05 1.2e-05
Small-block Lanczos Block Lanczos
Block size $d=8$ $d=16$ $d=32$ $d=64$ $d=128$ $d=128$
##\##matvec 1608 1632 1664 2048 3584 3584
Time (s) 48.55 43.52 40.29 44.42 70.72 69.70
$\lVert AV\rVert$ 1.0e-03 1.0e-03 1.0e-03 1.1e-03 1.0e-03 7.5e-04
5.3.3 Large-scale problems
In the final experiment, we address a particularly challenging problem where both the dense method and the traditional block Lanczos method are infeasible due to memory constraints. For the matrix “GL7d14”, it is impossible to apply the null function because the dense matrix cannot be stored in memory. Similarly, the large-block Lanczos method encounters a memory bottleneck, making it impractical to perform a restart after even just two iterations. Table 7 demonstrates that the the small-block Lanczos method successfully computes the null space within a Krylov subspace of dimension 10240. Remarkably, this dimension is only 1.29 times larger than the nullity, highlighting the efficiency and feasibility of the small-block method in handling such large-scale problems.
Table 7: Number of matrix vector multiplications and execution time of Algorithm 2 applied to “GL7d14” with nullity $7939$. The maximum dimension of Krylov subspace before restarting is $10240$. The computed nullity is correct and $\lVert AV\rVert\approx 1.4\times 10^{-4}$
Block size $d=16$ $d=32$ $d=64$ $d=128$ $d=256$ $d=512$
##\##matvec 133456 133344 132288 132352 133888 140800
Time(s) 11516 8137.4 6986.4 6375.0 6218.7 6716.4
6 Conclusion
In this work, we proposed a novel randomized small-block Lanczos method that incorporates modern Krylov subspace techniques such as restarting and partial reorthogonalization to compute null spaces of large sparse matrices. For the examples considered, our implementation outperforms both the default SVD-based null space solver and the traditional block Lanczos method. Additionally, we provide a theoretical analysis of the convergence for the single-vector case, that is, $d=1$. It remains an open problem to derive convergence bounds that reflect the good performance observed for small $d>1$.
Acknowledgments
The authors thank Yuji Nakatsukasa and Rikhav Shah for helpful discussions on this work.